Solve the following system by substitution
8x-4y=16
y=2x-4

Answer by jainenderkapoor(61)

8x-4y=16 ----(1)
y = 2x - 4 ----(2)
From the second equation
we have
y =2x - 4 ----- (3)
substituting this value in (1) equation, we get
8x - 3(2x - 4) = 16
8x - 6x + 12 = 16
2x + 12 = 16
2x + 12 - 12 = 16 - 12 (subtractinc 12 from both sides)
2x = 4
2x/2 = 4/2 (dividing both sides by 2)
x = 2
Putting this value of x in (3) equation we get
y = 2(2) - 4
y = 4 - 4
y = 0
so solution is x = 2 and y = 0
Hope the solution is clear to you. In case you have any other doubt, please contact me on jainenderkapoor@gmail.com I am available for Online Tutoring at very reasonable rates.